# A Probabilistic Approach to Uncertainty Quantification Leveraging 3D Geometry

### v1

@@ 5.3 Calculating the Hessian
The uncertainty computation in BayesSDF revolves around accumulating partial derivatives of the rendered color $\mathbf{c}(\mathbf{x})=(r(\mathbf{x}),\,g(\mathbf{x}),\,b(\mathbf{x}))$ with respect to the deformation parameters stored in the hash grid. Because the deformation field is queried at multiple corners (8 corners per voxel), we track partial derivatives for each corner contribution. Let us denote the final, aggregated color at a particular sample by $\mathbf{C}\;=\;\sum_{\mathbf{x}\in\mathcal{R}\text{-samples}}\mathbf{c}(\mathbf{x})$ where the sum is over all the sample points along the ray (or potentially aggregated over multiple rays). In the code, we backpropagate through each color channel independently to accumulate partial derivatives. For a single color channel $c\in\{r,g,b\}$, we compute $\frac{\partial c}{\partial\mathbf{d}_{c}^{(k)}}$ where $\mathbf{d}_{c}^{(k)}$ refers to the offsets (the deformation parameters) at corner $k\in\{1,\dots,8\}$. We do this by invoking a backwards pass for each color channel and capturing the resulting gradients with respect to $\mathbf{d}$. Each time, we zero out the gradient from the previous backward pass to isolate the new color channel’s derivative. Between each step, the accumulated gradient is saved (and subsequently multiplied by interpolation coefficients, etc.), ensuring that the next backward pass only captures derivatives for the next channel. This pattern also addresses memory constraints: if we did not zero out or separate channel-wise backprop, the gradient buffers for the entire 3D field could become unwieldy. Once these partial derivatives are obtained, we aggregate them in a variable named hessian in the script. Concretely, for each corner’s index k in the hash grid, the squared sum of partial derivatives is computed as follows:
@@
where $\mathbf{d}_{d}\in\{\mathbf{d}_{1},\mathbf{d}_{2},\mathbf{d}_{3}\}$ is each dimension of the deformation vector. In short, we take the partial derivative of r, g, and b with respect to each dimension of the local offset, square them, and sum them to form a scalar representing local “curvature” in color space for that grid corner. We organize these per-corner accumulations into a 3D tensor of size $((2^{\ell})+1)^{3}$. Each cell in that 3D tensor corresponds to a location in the hash-encoded grid. Over multiple training iterations, we repeatedly add these partial-derivative-based values, accumulating them into a global measure that, once normalized or post-processed, can be interpreted as the Hessian with respect to the neural color output and the deformation parameters.
5.4 Perturbations in Distortion Space
A critical aspect of BayesSDF is its reliance on small local changes in the deformation field to gauge uncertainty. Formally, if $\mathbf{d}(\mathbf{x})$ is the deformation vector for the voxel corner associated with point $\mathbf{x}$, we are implicitly examining the infinitesimal perturbations $\delta\mathbf{d}(\mathbf{x})\;\in\;\mathbb{R}^{3}$ and measuring how they affect the rendered color $\mathbf{c}(\mathbf{x})$. In other words, we approximate the local mapping
@@ 5.4 Perturbations in Distortion Space
where $\nabla_{\mathbf{d}(\mathbf{x})}\mathbf{c}(\mathbf{x})$ is precisely the gradient tensor we accumulate in the Hessian array. The magnitude of these partial derivatives thus quantifies the sensitivity of $\mathbf{c}(\mathbf{x})$ to local perturbations $\delta\mathbf{d}(\mathbf{x})$. In practice, perturbations are never explicitly applied to $\mathbf{d}(\mathbf{x})$; rather, their effect is simulated by the gradient computation (.backward()). This is central to the BayesSDF mechanism: we do not need to stochastically perturb each voxel corner. Instead, automatic differentiation yields $\partial\mathbf{c}(\mathbf{x})/\partial\mathbf{d}(\mathbf{x})$ for each channel, which we then square and sum to approximate the local uncertainty.
5.5 Approximating Uncertainty
We interpret these aggregated squared gradients in hessian as proxies for uncertainty. High squared partial derivatives indicate that small changes in the deformation field cause large changes to the rendered color, suggesting the reconstruction at that location is sensitive—and potentially uncertain. Conversely, near-zero partial derivatives indicate that changes in local deformation parameters barely affect the color, suggesting more stable or confident estimates. Given the final accumulated $\mathbf{H}$ stored in our 3D grid, we define an uncertainty function $\sigma$ over the domain:
@@ 5.5 Approximating Uncertainty
Because $\mathbf{H}$ is computed by summing the squared derivatives for the three color channels (each with respect to three offset dimensions), it encapsulates the total “sensitivity” at a location. One may further process $\sigma(\mathbf{x})$ by normalization or thresholding to highlight regions of highest uncertainty.
6 Experiments
@@ 6.1 Rendering Techniques
We employ a custom rendering method that integrates seamlessly with BayesSDF to generate RGB, depth, and uncertainty images under various camera trajectories. This rendering process is a vital component of our experimental pipeline, enabling us to visualize and analyze the uncertainty predictions alongside the scene geometry. The method supports producing multiple outputs, most importantly: RGB images, Depth images (visualized with a colormap), and Uncertainty maps (also colormapped, typically “inferno”). Each pixel’s uncertainty value is fed through a user-specified color mapping function to produce interpretive 2D heatmaps of the 3D uncertainty field.
Figure 1: Africa Dataset BayesSDF Uncertainty Rendering (Left: RGB, Middle: Uncertainty, Right: Depth)
Figure 2: Basket Dataset BayesSDF Uncertainty Rendering (Left: Depth, Middle: Uncertainty, Right: RGB)
@@ 6.2 Evaluation Metrics
Figure 3: Africa Dataset $\Delta\text{MAE}$ Sparsification Curve
Figure 4: Basket Dataset $\Delta\text{MAE}$ Sparsification Curve
Figure 5: Africa Dataset Ensemble Vs. BayesSDF Method Comparison
Figure 6: Basket Dataset Ensemble Vs. BayesSDF Method Comparison
@@
    where $\hat{d}(u,v)$ is the predicted depth, $d^{(}u,v)$ is ground-truth depth, and $\Omega(p)$ is the set of remaining pixels after removing a fraction p of the domain based on uncertainty sorting. Plotting $\Delta\text{MAE}$ against the fraction of pixels removed quantifies how well our uncertainty measure correlates with actual reconstruction errors. If the model is correct that “most uncertain” pixels also have high error, $\Delta\text{MAE}$ decreases rapidly when those pixels are removed.
  • Area Under the Sparsification Error (AUSE) with MAE: We integrate $\Delta\text{MAE}$ across the entire sparsification curve. A lower AUSE value indicates better alignment between “predicted uncertainty” and “true error,” meaning the model is effectively ranking pixels by their likelihood of error.
@@ 6.3 Results and Analysis
Table 2: Area Under the Sparsification Error (AUSE) with Mean Absolute Error (MAE) across different datasets and methods. Lower values indicate better performance.
@@
    Single-Pass Hessian Approximation: Rather than training multiple models (ensembles) or repeatedly sampling dropouts, BayesSDF captures uncertainty by analyzing local color-space derivatives with respect to the scene’s underlying deformation field. This approach is both memory and compute efficient.
@@
    Geometry-Driven Uncertainty: By tying uncertainty with signed distance function’s sensitivity, BayesSDF provides a physically interpretable metric. Regions where tiny deformation changes provoke large color changes correspond exactly to geometry that is poorly constrained or ambiguous.
@@
To validate that BayesSDF’s uncertainty maps correspond to poor geometry, we compare high-uncertainty regions against ground-truth geometry across all datasets. We observe that high Hessian-based uncertainty consistently aligns with areas exhibiting large depth errors or missing surface detail such as reflective regions, occlusions, or sparse view coverage. These findings suggest that our Laplace-style approximation effectively identifies instability in the underlying SDF geometry, making the uncertainty not just a confidence metric but a reliable indicator of physical reconstruction error.
@@ 7 Applications
The BayesSDF framework offers a transformative approach to uncertainty quantification, with significant implications for several domains that require precise and reliable 3D reconstructions. Below, we discuss the core applications of BayesSDF, emphasizing its technical strengths and relevance across various fields. BayesSDF is particularly well-suited for dense 3D scene reconstruction, a critical task in computer vision and robotics. By leveraging ensemble-based uncertainty, BayesSDF quantifies spatial ambiguities in reconstructed regions. These uncertainty maps enable robotic systems to detect areas of high ambiguity, which is invaluable for autonomous navigation in dynamic or cluttered environments. For example, in Simultaneous Localization and Mapping (SLAM) systems, BayesSDF can improve loop closure detection by flagging uncertain regions in the map that require additional measurements. Physics-based simulations in fields such as engineering and medical imaging rely heavily on accurate 3D models to simulate phenomena such as fluid dynamics or stress distribution. BayesSDF enhances these simulations by providing uncertainty-aware geometric models. For instance, in medical imaging, BayesSDF could quantify the variability in 3D reconstructions of anatomical structures, such as bones or blood vessels, to assess confidence in diagnostic or treatment planning tools. BayesSDF opens new avenues in scientific research fields that depend on precise 3D modeling, such as computational biology and geology. For example, in paleontology, BayesSDF can assist in reconstructing fossilized remains with quantified confidence levels, allowing researchers to distinguish between highly reliable and uncertain reconstructions.
@@ 8 Conclusion
We introduce BayesSDF, a probabilistic framework designed for uncertainty quantification in SDF-based 3D reconstruction models. Motivated by scientific and physical simulation applications where accurate surface geometry is paramount, BayesSDF integrates a Laplace approximation technique over signed distance fields to quantify local reconstruction instability. Prior work has established that SDFs yield smoother, more consistent surfaces than NeRFs, and our method builds on this advantage to provide meaningful geometric uncertainty. The result is a compact, computationally efficient uncertainty estimator that aligns uncertainty directly with surface quality. By utilizing models like NeuS’ volumetric SDF space, BayesSDF provides uncertainty maps that are inherently aligned with the reconstructed surfaces, enabling reliable and precise uncertainty quantification. The technical contributions of BayesSDF include a deformation field model based on hierarchical hashing, a computationally efficient Hessian-based curvature estimation, and ensemble-based epistemic uncertainty metrics. Together, these components address critical challenges in 3D uncertainty estimation, including the misalignment of uncertainty with surface geometry and the high computational costs associated with traditional sampling-based approaches. Through rigorous experiments on synthetic and real-world datasets, BayesSDF demonstrates outstanding performance, achieving very high calibration results and reduced computational overhead. Looking forward, the principles of BayesSDF can be extended to broader applications, including dynamic scene understanding and hybrid models that combine SDFs with density-based neural representations. Moreover, the integration of BayesSDF with hardware-accelerated systems and distributed processing frameworks presents exciting opportunities for real-time 3D uncertainty quantification.

### v2

@@ 5.3 Calculating the Hessian
The uncertainty computation in BayesSDF revolves around accumulating partial derivatives of the rendered color $\mathbf{c}(\mathbf{x})=(r(\mathbf{x}),\,g(\mathbf{x}),\,b(\mathbf{x}))$ with respect to the deformation parameters stored in the hash grid. Because the deformation field is queried at multiple corners (8 corners per voxel), we track partial derivatives for each corner contribution. Let us denote the final, aggregated color at a particular sample by $\mathbf{C}\;=\;\sum_{\mathbf{x}\in\mathcal{R}\text{-samples}}\mathbf{c}(\mathbf{x})$ where the sum is over all the sample points along the ray (or potentially aggregated over multiple rays). In the code, we backpropagate through each color channel independently to accumulate partial derivatives. For a single color channel $c\in\{r,g,b\}$, we compute $\frac{\partial c}{\partial\mathbf{d}_{c}^{(k)}}$ where $\mathbf{d}_{c}^{(k)}$ refers to the offsets (the deformation parameters) at corner $k\in\{1,\dots,8\}$. We do this by invoking a backwards pass for each color channel and capturing the resulting gradients with respect to $\mathbf{d}$. Each time, we zero out the gradient from the previous backward pass to isolate the new color channel’s derivative. Between each step, the accumulated gradient is saved (and subsequently multiplied by interpolation coefficients, etc.), ensuring that the next backward pass only captures derivatives for the next channel. This pattern also addresses memory constraints: if we did not zero out or separate channel-wise backprop, the gradient buffers for the entire 3D field could become unwieldy. Once these partial derivatives are obtained, we aggregate them in a variable called ‘hessian‘ in the script. Concretely, for each corner’s index k in the hash grid, the squared sum of partial derivatives is computed as follows:
@@
where $\mathbf{d}_{d}\in\{\mathbf{d}_{1},\mathbf{d}_{2},\mathbf{d}_{3}\}$ represents each dimension of the deformation vector. In short, we take the partial derivative of r, g, and b with respect to each dimension of the local offset, square them, and sum them to form a scalar representing local “curvature” in color space for that grid corner. We organize these per-corner accumulations into a 3D tensor of size $((2^{\ell})+1)^{3}$. Each cell in that 3D tensor corresponds to a location in the hash-encoded deformation grid. Over multiple training iterations, we repeatedly add these partial-derivative-based values, accumulating them into a global measure that, once normalized or post-processed, can be interpreted as the Hessian with respect to the neural color output and the deformation parameters.
5.4 Distortion Space Perturbations
A critical aspect of BayesSDF is its reliance on small local changes in the deformation field to gauge uncertainty. Formally, if $\mathbf{d}(\mathbf{x})$ is the deformation vector for the voxel corner associated with point $\mathbf{x}$, we implicitly examine the infinitesimal perturbations $\delta\mathbf{d}(\mathbf{x})\;\in\;\mathbb{R}^{3}$ measure their effect on the rendered color $\mathbf{c}(\mathbf{x})$. In other words, we approximate the local mapping
@@ 5.4 Distortion Space Perturbations
where $\nabla_{\mathbf{d}(\mathbf{x})}\mathbf{c}(\mathbf{x})$ is precisely the gradient tensor we accumulate in the Hessian array. The magnitude of these partial derivatives quantifies the sensitivity of $\mathbf{c}(\mathbf{x})$ to local perturbations $\delta\mathbf{d}(\mathbf{x})$. In practice, perturbations are never explicitly applied to $\mathbf{d}(\mathbf{x})$; rather, their effect is simulated by the gradient computation (.backward()). This is central to the BayesSDF mechanism: we do not need to stochastically perturb each voxel corner. Instead, automatic differentiation yields $\partial\mathbf{c}(\mathbf{x})/\partial\mathbf{d}(\mathbf{x})$ for each channel, which we then square and sum to approximate local uncertainty.
5.5 Uncertainty Approximation
We interpret these aggregated squared gradients in ‘hessian‘ as proxies for uncertainty. High squared partial derivatives indicate that small changes in the deformation field cause large changes to the rendered color, suggesting the reconstruction at that location is sensitive—and potentially uncertain. Conversely, near-zero partial derivatives indicate that changes in local deformation parameters barely affect the color, suggesting more stable or confident estimates. Given the final accumulated $\mathbf{H}$ stored in our 3D grid, we define an uncertainty function $\sigma$ over the domain:
@@ 5.5 Uncertainty Approximation
Because $\mathbf{H}$ is computed by adding the squared derivatives for the three color channels (each with respect to three offset dimensions), it encapsulates the total ’sensitivity’ at a location. One may further process $\sigma(\mathbf{x})$ by normalization or thresholding to highlight regions of highest uncertainty.
6 Results
@@ 6.1 Rendering Techniques
We employ a custom rendering method that integrates seamlessly with BayesSDF to generate RGB, depth, and uncertainty images under various camera trajectories. This rendering process is a vital component of our experimental pipeline, enabling us to visualize and analyze uncertainty predictions alongside scene geometry. The method supports producing multiple outputs, most importantly: RGB images, Depth images (visualized with a colormap), and Uncertainty maps (also colormapped, typically “inferno”). Each pixel’s uncertainty value is fed through a user-specified color mapping function to produce interpretive 2D heatmaps of the 3D uncertainty field.
Figure 1: Africa Dataset: BayesSDF-derived view with uncertainty rendering. The left panel shows the RGB image, the middle panel depicts predictive uncertainty, and the right panel illustrates depth.
Figure 2: Basket Dataset: BayesSDF-generated view with uncertainty. The left panel displays the depth map, the middle panel shows the uncertainty estimation, and the right panel presents the RGB image.
@@ 6.2 Evaluation Metrics
Figure 3: Africa Dataset: Change in mean absolute error ($\Delta$MAE) as a function of sparsification.
Figure 4: Basket Dataset: Sparsification curve showing the change in mean absolute error ($\Delta$MAE) with increasing levels of sparsification.
Figure 5: Africa Dataset: Comparison of the Ensemble method and BayesSDF in terms of AUSE.
Figure 6: Basket Dataset: Comparison of the Ensemble method and BayesSDF in terms of AUSE.
@@
    where $\hat{d}(u,v)$ is the predicted depth, $d^{(}u,v)$ is ground-truth depth, and $\Omega(p)$ is the set of remaining pixels after removing a fraction p of the domain based on uncertainty sorting. Plotting $\Delta\text{MAE}$ against the fraction of pixels removed quantifies how well our uncertainty measure correlates with actual reconstruction errors. If the model correctly identifies ”most uncertain” pixels as having high error, $\Delta\text{MAE}$ decreases rapidly when those pixels are removed.
  • Area Under the Sparsification Error (AUSE) with MAE: We integrate $\Delta\text{MAE}$ over the entire sparsification curve. A lower AUSE value indicates better alignment between “predicted uncertainty” and “true error,” indicating the model is effectively ranking pixels by their likelihood of error.
@@ 6.3 Analysis
Table 2: Area Under the Sparsification Error (AUSE) with Mean Absolute Error (MAE) for various datasets and reconstruction methods. Lower AUSE values indicate superior performance.
@@
    Single-Pass Hessian Approximation: Rather than training multiple models (ensembles) or repeatedly sampling dropouts, BayesSDF captures uncertainty by analyzing local colorspace derivatives with respect to the scene’s underlying deformation field. This approach is both memory and compute efficient.
@@
    Geometry-Driven Uncertainty: BBy relating uncertainty to the signed distance function’s sensitivity, BayesSDF provides a physically interpretable metric. Regions where small deformation changes cause large color changes correspond to geometry that is poorly constrained or ambiguous.
@@
To validate that BayesSDF’s uncertainty maps correspond to poor geometry, we compare high-uncertainty regions against ground-truth geometry across all datasets. We observe that high Hessian-based uncertainty consistently aligns with areas exhibiting large depth errors or missing surface detail such as reflective regions, occlusions, or sparse view coverage. These findings suggest that our Laplace approximation effectively identifies instability in the underlying SDF geometry, making the uncertainty not just a confidence metric but a reliable indicator of physical reconstruction error.
@@ 7 Applications
The BayesSDF framework offers a transformative approach to uncertainty quantification, with significant implications for several domains that require precise and reliable 3D reconstructions. Below, we discuss the core applications of BayesSDF, emphasizing its technical strengths and relevance across various fields. BayesSDF is particularly well-suited for dense 3D scene reconstruction, a critical task in computer vision and robotics. BayesSDF quantifies spatial ambiguities in reconstructed regions. These uncertainty maps enable robotic systems to detect areas of high ambiguity, invaluable for autonomous navigation in dynamic or cluttered environments. For example, in Simultaneous Localization and Mapping (SLAM) systems, BayesSDF can improve loop closure detection by flagging uncertain regions in the map that require additional measurements. Physics-based simulations in fields such as engineering and medical imaging rely on accurate 3D models to simulate phenomena such as fluid dynamics or stress distribution. BayesSDF enhances these simulations by providing geometric models with quantified uncertainty. For instance, in medical imaging, BayesSDF could quantify the variability in 3D reconstructions of anatomical structures, such as bones or blood vessels, to assess confidence in diagnostic or treatment planning tools. BayesSDF opens new avenues in scientific research fields that depend on precise 3D modeling, such as computational biology and geology. For example, in paleontology, BayesSDF can assist in reconstructing fossilized remains with quantified confidence levels, allowing researchers to distinguish between highly reliable and uncertain reconstructions.
@@ 8 Conclusion
We introduce BayesSDF, a probabilistic framework designed for uncertainty quantification in SDF-based 3D reconstruction models. Motivated by scientific and physical simulation applications where accurate surface geometry is paramount, BayesSDF integrates a Laplace approximation technique over signed distance fields to quantify local reconstruction instability. Prior work has established that SDFs yield smoother, more consistent surfaces than NeRFs, and our method builds on this advantage to provide meaningful geometric uncertainty. The result is a compact, efficient uncertainty estimator that aligns uncertainty directly with surface quality. By utilizing models like NeuS’ volumetric SDF space, BayesSDF provides uncertainty maps that are inherently aligned with the reconstructed surfaces, enabling reliable and precise uncertainty quantification. The technical contributions of BayesSDF include a deformation field model based on hierarchical hashing, a computationally efficient Hessian-based curvature estimation, and ensemble-based epistemic uncertainty metrics. Together, these components address critical challenges in 3D uncertainty estimation, including the misalignment of uncertainty with surface geometry and the high computational costs associated with traditional sampling-based approaches. Through rigorous experiments on synthetic and real-world datasets, BayesSDF demonstrates outstanding performance, achieving very high calibration results and reduced computational overhead. Looking forward, the principles of BayesSDF can be extended to broader applications, such as dynamic scene understanding and hybrid models that combine SDFs with density-based neural representations. Moreover, the integration of BayesSDF with hardware-accelerated systems and distributed processing frameworks presents exciting opportunities for real-time 3D uncertainty quantification.